# Long-Context Linear System Identification

### v1

@@ B.2 Technical lemmas
$\begin{split}{{\left\|\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right\|}}_{2}^{2}&\leqslant 2{{\left\|\left(I-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}\right\|}}_{\rm op}\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right)\\ &=2{{\left\|I_{Td}+\left(M_{{\boldsymbol{A}}^{\star}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}\right\|}}_{\rm op}\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right)\\ &\leqslant 2\left(1+{{\left\|\left(M_{{\boldsymbol{A}}^{\star}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}\right\|}}_{\rm op}\right)\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right)\\ &=2\eta\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right)\,.\end{split}$
B.3 Lower and upper isometries
In Theorem B.16, we present a uniform lower bound on $\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}$. In order to establish this lower bound, we first start with a point-wise lower bounds in Lemma B.19 that relies on Hanson-Wright inequality for bounding the deviation of quadratic forms of sub-Gaussian vectors. Then, we use Lemmas B.21 and B.22 with a discretization argument in Theorem B.16 to establish uniform isometries. Finally, with Corollary B.17, we have a uniform control over the range of both $\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}$ and $\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)$.
Theorem B.16.
Let $\delta>0$ be small and fixed. Then, there exist a constant $1\leqslant C(\delta)=\ln(\frac{1}{\delta})$ such that the following holds uniformly for all ${\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)$ and $C\geqslant C(\delta)$:
$\begin{split}\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}&\geqslant\frac{\sigma^{2}}{8}\sigma_{\min}(L_{\star})^{2}N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}\,,\\ \end{split}$ (26)
@@ B.3 Lower and upper isometries
By Lemmas B.20 and B.19, with probability at least $1-\delta_{1}-\delta_{2}$, the following holds:
$\begin{split}\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}&\geqslant\sigma^{2}\left(1-c^{2}\nu_{1}\right)\sigma_{\min}(L_{\star})^{2}N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}\,,\\ \end{split}$ (27)
for any arbitrary ${\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)$ where
$\delta_{1}=\exp\left(-C_{HW}C\nu_{1}^{2}p^{\prime}dr\ln\tau\right)\,,\quad\delta_{2}=\exp\left(-C_{HW}C\nu_{2}^{2}p^{\prime}dr\ln\tau\right)\,.$
@@
the condition Equation 27 holds $\forall{\boldsymbol{A}}\in\mathcal{N}_{\epsilon}(D)$. Moreover, by Lemmas B.21, B.22 and B.23, we have
$\begin{split}\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}&\geqslant\frac{1}{2}\sigma^{2}\left(1-c^{2}\nu_{1}\right)\sigma_{\min}(L_{\star})^{2}N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}-\sigma^{2}(1+c^{2}\nu_{3})\epsilon^{2}\|L_{\star}\|_{\rm op}^{2}p^{\prime}dNT\,\\ \operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)&\leqslant\sigma^{2}c^{2}\nu_{2}\|L_{\star}\|_{\rm op}\sqrt{Cp^{\prime}drN\ln\tau}\|\Delta_{\boldsymbol{A}}\|_{F}+\sigma^{2}(1+c^{2}\nu_{3})\epsilon\|L_{\star}\|_{\rm op}\sqrt{p^{\prime}}dNT\,,\end{split}$
@@
$\delta_{3}=\exp\left(-C_{HW}\nu_{3}dNT\right),\,.$
@@
Setting $\nu_{1}=\frac{1}{4c^{2}}$, $\nu_{2}=\frac{1}{4c^{2}}$ and $\epsilon$ such that
$\epsilon=\frac{1}{2\left(1+c^{2}\nu_{3}\right)}\cdot\sqrt{\dfrac{T-p^{\prime}}{T}}\min{{\left\{\frac{1}{\sigma_{\mathrm{cond}}(L_{\star})}\sqrt{\dfrac{r}{dNT}},\sqrt{\dfrac{1}{p^{\prime}d}}\right\}}}\,,$
and recalling that $C\geqslant 1$ and $\ln\tau\geqslant 1$,
$\begin{split}\frac{1}{2}\sigma^{2}\left(1-c^{2}\nu_{1}\right)\sigma_{\min}(L_{\star})^{2}N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}&-\sigma^{2}(1+c^{2}\nu_{3})\epsilon^{2}\|L_{\star}\|_{\rm op}^{2}p^{\prime}dNT\geqslant\frac{\sigma^{2}}{8}\sigma_{\min}(L_{\star})^{2}N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}\,,\\ \sigma^{2}c^{2}\nu_{2}\|L_{\star}\|_{\rm op}\sqrt{Cp^{\prime}drN\ln\tau}\|\Delta_{\boldsymbol{A}}\|_{F}&+\sigma^{2}(1+c^{2}\nu_{3})\epsilon\|L_{\star}\|_{\rm op}\sqrt{p^{\prime}}dNT\\ \quad&\leqslant\frac{\sigma^{2}}{2}\|L_{\star}\|_{\rm op}\sqrt{Cp^{\prime}drN\ln\tau}\|\Delta_{\boldsymbol{A}}\|_{F}\,,\end{split}$
@@
$\begin{split}\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}&\geqslant\frac{\sigma^{2}}{8}\sigma_{\min}(L_{\star})^{2}N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}\,,\\ \operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)&\leqslant\frac{\sigma^{2}}{2}\|L_{\star}\|_{\rm op}\sqrt{Cp^{\prime}drN\ln\tau}\|\Delta_{\boldsymbol{A}}\|_{F}\,.\end{split}$
Lastly, we can ensure that $\delta_{3}<\delta/2$ with the choice of
$\nu_{3}(\delta)=\dfrac{\ln\frac{1}{\delta/2}}{C_{HW}}>\dfrac{\ln\frac{1}{\delta/2}}{C_{HW}dNT}\,.$
Moreover, the $\epsilon$-net size can be bounded as follows:
$|\mathcal{N}_{\varepsilon}(D)|\leqslant\left(\frac{9}{\epsilon}\right)^{(p^{\prime}d+d+1)r}\leqslant\exp\left(3p^{\prime}dr\ln\frac{9}{\epsilon}\right)\leqslant\exp\left(9p^{\prime}dr\ln\frac{1}{\epsilon}\right)\,.$
For more details on $\epsilon$-nets on low-rank martices, see Candès & Plan (2011, Lemma 3.1). Then,
$\delta_{0}=|\mathcal{N}_{\epsilon}(D)|\left(\delta_{1}+\delta_{2}\right)\leqslant\exp\left(\left(9-\frac{1}{16c^{4}}C_{HW}C\right)p^{\prime}dr\ln\tau+9p^{\prime}dr\ln\left(1+\dfrac{c^{2}}{C_{HW}}\ln\frac{1}{\delta/2}\right)\right)\,,$
where we use that $\frac{1}{\epsilon}<\tau\left(1+\dfrac{c^{2}}{C_{HW}}\ln\frac{1}{\delta/2}\right)$. Thus, $\delta_{0}<\delta/2$ can be made with the choice of
$\begin{split}C(\delta)&=\frac{16c^{4}}{C_{HW}}\left(9+\ln\frac{1}{\delta/2}+9\ln\left(1+\dfrac{c^{2}}{C_{HW}}\ln\frac{1}{\delta/2}\right)\right)\\ &>\frac{16c^{4}}{C_{HW}}\left(9+\dfrac{\ln\frac{1}{\delta/2}}{p^{\prime}dr\ln\tau}+9\dfrac{\ln\left(1+\dfrac{c^{2}}{C_{HW}}\ln\frac{1}{\delta/2}\right)}{\ln\tau}\right)\,.\end{split}$
Here, $C(\delta)$ is a constant that is independent of $p,p^{\prime},d,r,N,T$ such that $C(\delta)=\mathcal{O}(\ln(1/\delta))$. ∎
Corollary B.17.
For any small $\delta>0$, there exists a constant $1\leqslant C(\delta)=\mathcal{O}(\ln(1/\delta))$ such that the following holds uniformly for all ${\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)$ and $C\geqslant C(\delta)$:
$\begin{split}\inf_{{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)}\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}&\geqslant\frac{\sigma^{2}}{8}\sigma_{\min}(L_{\star})^{2}C\eta^{2}D^{2}p^{\prime}dr\left(\ln\tau\right)^{2}\,,\\ \sup_{{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)}\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)&\leqslant\frac{\sigma^{2}}{2}\|L_{\star}\|_{\rm op}C\eta Dp^{\prime}dr\left(\ln\tau\right)^{3/2}\,,\end{split}$ (28)
@@
Plug in the results from Theorem B.16 and use the definition of the set $\mathcal{A}(D)\setminus\mathcal{G}(C,D)$. ∎
Definition B.18.
For applying Theorem A.1 in our setup, consider the following objects:
@@
$1-\exp\left(-C_{HW}C\nu^{2}p^{\prime}dr\ln\tau\right)\,,$
@@
Applying Theorem A.1 with $P=\tilde{\Delta}_{\boldsymbol{A}}^{\top}\tilde{\Delta}_{\boldsymbol{A}}$ and $r=c^{2}\sigma^{2}\nu\|\tilde{\Delta}_{\boldsymbol{A}}\|_{F}^{2}$,
$\begin{split}\mathbb{P}\big{(}\tilde{E}^{\top}\tilde{\Delta}_{\boldsymbol{A}}^{\top}\tilde{\Delta}_{\boldsymbol{A}}\tilde{E}-\mathbb{E}[\tilde{E}^{\top}\tilde{\Delta}_{\boldsymbol{A}}^{\top}\tilde{\Delta}_{\boldsymbol{A}}\tilde{E}]&\geqslant c^{2}\sigma^{2}\nu\|\tilde{\Delta}_{\boldsymbol{A}}\|_{F}^{2}\big{)}\\ \quad&\leqslant\exp\left(-C_{HW}\min{{\left\{\nu^{2}\frac{\|\tilde{\Delta}_{\boldsymbol{A}}\|_{F}^{4}}{\|\tilde{\Delta}_{\boldsymbol{A}}^{\top}\tilde{\Delta}_{\boldsymbol{A}}\|_{F}^{2}},\nu\frac{\|\tilde{\Delta}_{\boldsymbol{A}}\|_{F}^{2}}{\|\tilde{\Delta}_{\boldsymbol{A}}^{\top}\tilde{\Delta}_{\boldsymbol{A}}\|_{\rm op}}\right\}}}\right).\end{split}$
@@
$\mathbb{P}\left(\|\Delta_{\boldsymbol{A}}E\|_{F}^{2}-\sigma^{2}N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}\geqslant c^{2}\sigma^{2}\nu N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}\right)\leqslant\exp\left(-C_{HW}\min{{\left\{\nu^{2}N\frac{\|\Delta_{\boldsymbol{A}}\|_{F}^{4}}{\|\Delta_{\boldsymbol{A}}^{\top}\Delta_{\boldsymbol{A}}\|_{F}^{2}},\nu N\frac{\|\Delta_{\boldsymbol{A}}\|_{F}^{2}}{\|\Delta_{\boldsymbol{A}}\|_{\rm op}^{2}}\right\}}}\right).$
@@
$\begin{split}1-\exp\left(-C_{HW}C\nu^{2}p^{\prime}dr\ln\tau\right)\,,\end{split}$
@@
$\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)\leqslant c^{2}\sigma^{2}\nu\|L_{\star}\|_{\rm op}\sqrt{Cp^{\prime}drN\ln\tau}\|\Delta_{\boldsymbol{A}}\|_{F}\,.$
@@
Applying Theorem A.1 with $P=\tilde{\Delta}_{\boldsymbol{A}}$ and $r=c^{2}\sigma^{2}\nu\sqrt{Cp^{\prime}dr\ln\tau}\|\tilde{\Delta}_{\boldsymbol{A}}\|_{F}$,
$\begin{split}&\mathbb{P}\left(\tilde{E}^{\top}\tilde{\Delta}_{\boldsymbol{A}}\tilde{E}-\mathbb{E}[\tilde{E}^{\top}\tilde{\Delta}_{\boldsymbol{A}}\tilde{E}]\geqslant c^{2}\sigma^{2}\nu\sqrt{Cp^{\prime}dr\ln\tau}\|\tilde{\Delta}_{\boldsymbol{A}}\|_{F}\right)\\ &\quad\quad\leqslant\exp\left(-C_{HW}\min{{\left\{\nu^{2}Cp^{\prime}dr\ln\tau,\nu\sqrt{Cp^{\prime}dr\ln\tau}\frac{\|\tilde{\Delta}_{\boldsymbol{A}}\|_{F}}{\|\tilde{\Delta}_{\boldsymbol{A}}\|_{\rm op}}\right\}}}\right).\end{split}$
@@
$\begin{split}&\mathbb{P}\left(\operatorname{Tr}\left(E^{\top}\Delta_{\boldsymbol{A}}E\right)\leqslant c^{2}\sigma^{2}\nu\sqrt{Cp^{\prime}drN\ln\tau}\|\Delta_{\boldsymbol{A}}\|_{F}\right)\geqslant\\ &\qquad\qquad\qquad 1-\exp\left(-C_{HW}\min{{\left\{\nu^{2}Cp^{\prime}dr\ln\tau,\nu\sqrt{Cp^{\prime}drN\ln\tau}\frac{\|\Delta_{\boldsymbol{A}}\|_{F}}{\|\Delta_{\boldsymbol{A}}\|_{\rm op}}\right\}}}\right).\end{split}$
@@
Lemma B.21.
For any $\nu\geqslant 1$, with probability at least
$1-\exp\left(-C_{HW}\nu dNT\right),$
@@
Applying Theorem A.1 with $P=I_{dTN},r=c^{2}\sigma^{2}\nu dNT$,
@@
The result follows after a simple computation. ∎
Lemma B.22.
@@
$\begin{split}\|\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E\|_{F}^{2}&\geqslant\frac{1}{2}\|\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E\|_{F}^{2}-p^{\prime}\|{\boldsymbol{A}}_{1}-{\boldsymbol{A}}_{2}\|_{\rm op}^{2}\|L_{\star}\|_{\rm op}^{2}\|E\|_{F}^{2}\,.\end{split}$ (29)

### v2

@@ C.2 Technical lemmas
$\begin{split}{{\left\|\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right\|}}_{F}^{2}&\leqslant 2{{\left\|\left(I-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}\right\|}}_{\rm op}\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right)\\ &=2{{\left\|I_{Td}+\left(M_{{\boldsymbol{A}}^{\star}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}\right\|}}_{\rm op}\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right)\\ &\leqslant 2\left(1+{{\left\|\left(M_{{\boldsymbol{A}}^{\star}}-M_{{\boldsymbol{A}}^{\star}_{p^{\prime}}}\right)L_{\star}\right\|}}_{\rm op}\right)\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right)\\ &=2\eta\operatorname{Tr}\left(E^{\top}\Delta_{{\boldsymbol{\hat{A}}}}L_{\star}E\right)\,.\end{split}$
C.3 Lower and upper isometries
In Theorem C.3, we present uniform bounds on $\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}$ and $\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)$ in terms of $\|\Delta_{\boldsymbol{A}}\|_{F}$. In order to establish these bounds, we first start with point-wise bounds in Lemmas C.5 and C.6 that rely on Hanson-Wright inequality for bounding the deviation of quadratic forms. Then, we use Lemmas C.7, C.8 and C.9 with a discretization argument to establish uniform isometries. Finally, with Corollary C.5, we have a uniform control over the range of both $\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}$ and $\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)$.
Theorem C.3.
Let $\delta>0$ be small and fixed. Then, there exists a constant $1\leqslant C(\delta)=\mathcal{O}(\ln(1/\delta))$ such that the following holds uniformly for all ${\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)$ and $C\geqslant C(\delta)$:
$\begin{split}\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}&\geqslant\frac{\sigma^{2}}{8}\sigma_{\min}(L_{\star})^{2}N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}\,,\\ \operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)&\leqslant\sigma^{2}\|L_{\star}\|_{\rm op}\sqrt{C\tau p^{\prime}drN}\|\Delta_{\boldsymbol{A}}\|_{F}\,,\end{split}$
@@ C.3 Lower and upper isometries
Let $\nu_{1},\nu_{2}\in(0,1)$ be arbitrary. By Lemmas C.6 and C.5, with probability at least $1-\delta_{1}-\delta_{2}$, the following holds:
$\begin{split}\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}&\geqslant\sigma^{2}\left(1-c^{2}\nu_{1}\right)\sigma_{\min}(L_{\star})^{2}N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}\,,\\ \operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)&\leqslant\sigma^{2}c^{2}\nu_{2}\|L_{\star}\|_{\rm op}\sqrt{C\eta^{2}\tau p^{\prime}drN}\|\Delta_{\boldsymbol{A}}\|_{F}\,,\end{split}$ (23)
for any arbitrary $\nu_{1},\nu_{2}\in(0,1)$ and ${\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)$ where
$\delta_{1}=\exp\left(-C_{HW}C\nu_{1}^{2}\eta^{2}\tau p^{\prime}dr\right)\,,\quad\delta_{2}=\exp\left(-C_{HW}C\nu_{2}^{2}\eta^{2}\tau p^{\prime}dr\right)\,.$
@@
the condition Equation 23 holds $\forall{\boldsymbol{A}}\in\mathcal{N}_{\epsilon}(D)$. Moreover, by Lemmas C.7, C.8 and C.9, we have
$\begin{split}\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}&\geqslant\frac{1}{2}\sigma^{2}\left(1-c^{2}\nu_{1}\right)\sigma_{\min}(L_{\star})^{2}N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}-\sigma^{2}(1+c^{2}\nu_{3})\epsilon^{2}\|L_{\star}\|_{\rm op}^{2}p^{\prime}dNT\,\\ \operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)&\leqslant\sigma^{2}c^{2}\nu_{2}\|L_{\star}\|_{\rm op}\sqrt{C\eta^{2}\tau p^{\prime}drN}\|\Delta_{\boldsymbol{A}}\|_{F}+\sigma^{2}(1+c^{2}\nu_{3})\epsilon\|L_{\star}\|_{\rm op}\sqrt{p^{\prime}}dNT\,,\end{split}$
@@
$\delta_{3}=\exp\left(-C_{HW}\min{{\left\{\nu_{3},\nu_{3}^{2}\right\}}}dNT\right)\,.$
@@
Setting $\nu_{1}=\frac{1}{4c^{2}}$, $\nu_{2}=\frac{1}{2c^{2}}$ and $\epsilon$ such that
$\epsilon=\frac{1}{2\left(1+c^{2}\nu_{3}\right)}\sqrt{\dfrac{T-p^{\prime}}{T}}\min{{\left\{\frac{1}{\sigma_{\mathrm{cond}}(L_{\star})}\sqrt{\dfrac{1}{p^{\prime}d}},\sqrt{\dfrac{C\eta^{2}\tau r}{dNT}}\right\}}}\,,$
and we have the following:
$\begin{split}\frac{1}{2}\sigma^{2}\left(1-c^{2}\nu_{1}\right)\sigma_{\min}(L_{\star})^{2}N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}&-\sigma^{2}(1+c^{2}\nu_{3})\epsilon^{2}\|L_{\star}\|_{\rm op}^{2}p^{\prime}dNT\geqslant\frac{\sigma^{2}}{8}\sigma_{\min}(L_{\star})^{2}N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}\,,\\ \sigma^{2}c^{2}\nu_{2}\|L_{\star}\|_{\rm op}\sqrt{C\eta^{2}\tau p^{\prime}drN}\|\Delta_{\boldsymbol{A}}\|_{F}&+\sigma^{2}(1+c^{2}\nu_{3})\epsilon\|L_{\star}\|_{\rm op}\sqrt{p^{\prime}}dNT\\ \quad&\leqslant\sigma^{2}\|L_{\star}\|_{\rm op}\sqrt{C\eta^{2}\tau p^{\prime}drN}\|\Delta_{\boldsymbol{A}}\|_{F}\,,\end{split}$
@@
$\begin{split}\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}&\geqslant\frac{\sigma^{2}}{8}\sigma_{\min}(L_{\star})^{2}N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}\,,\\ \operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)&\leqslant\sigma^{2}\|L_{\star}\|_{\rm op}\sqrt{C\eta^{2}\tau p^{\prime}drN}\|\Delta_{\boldsymbol{A}}\|_{F}\,.\end{split}$
Lastly, we need ensure that $\delta_{0}+\delta_{3}\leqslant\delta$. First, $\delta_{3}\leqslant\delta/2$ can be achieved with the choice of
$\nu_{3}(\delta/2)=\max{{\left\{1,\dfrac{\ln\frac{1}{\delta/2}}{C_{HW}dNT}\right\}}}=\mathcal{O}\left(\ln(1/\delta)\right)\,.$
Moreover, the $\epsilon$-net size is bounded as follows:
$|\mathcal{N}_{\varepsilon}(D)|\leqslant\exp\left(9p^{\prime}dr\ln\frac{p^{\prime}}{\epsilon}\right)\,,$
by Corollary B.2. Then,
$\begin{split}\delta_{0}&=|\mathcal{N}_{\epsilon}(D)|\left(\delta_{1}+\delta_{2}\right)\\ &\leqslant\exp\left(-\frac{5}{16c^{4}}C_{HW}C\eta^{2}\tau p^{\prime}dr+9p^{\prime}dr\ln\frac{p^{\prime}}{\epsilon}\right)\,.\end{split}$
Thus, $\delta_{0}\leqslant\delta/2$ can be achieved with the choice of
$\begin{split}C\geqslant C(\delta)&=\frac{16c^{4}}{5C_{HW}\eta^{2}\tau}\left(9\ln\frac{p^{\prime}}{\epsilon}+\frac{\ln 2/\delta}{p^{\prime}dr}\right)\,.\\ \end{split}$ (24)
Note that $\eta,\tau\geqslant 1$ and the latter term is $\mathcal{O}(\ln(1/\delta))$. For the first term, crudely upper bound $\ln\frac{p^{\prime}}{\epsilon}$ by using $x\geqslant\ln x+1$ for any $x\geqslant 0$:
$\ln\frac{p^{\prime}}{\epsilon}\leqslant\ln\sqrt{\dfrac{p^{\prime 2}dNT^{2}}{T-p^{\prime}}}+\ln\sigma_{{\mathrm{cond}}}(L_{\star})+\ln 2(1+c^{2}\nu_{3}(\delta/2))\leqslant\tau+\mathcal{O}(\ln(1/\delta))\,.$
Therefore, the definition in Equation 24 verifies $C(\delta)=\mathcal{O}(\ln(1/\delta))$.
Corollary C.5.
For any small $\delta>0$, there exists a constant $1\leqslant C(\delta)=\mathcal{O}(\ln(1/\delta))$ such that the following holds uniformly for all ${\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C(\delta),D)$ and $C\geqslant C(\delta)$:
$\begin{split}\inf_{{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)}\|\Delta_{{\boldsymbol{A}}}L_{\star}E\|_{F}^{2}&\geqslant\frac{\sigma^{2}}{8}\sigma_{\min}(L_{\star})^{2}C(\delta)D^{2}\eta^{2}\tau p^{\prime}dr\,,\\ \sup_{{\boldsymbol{A}}\in\mathcal{A}(D)\setminus\mathcal{G}(C,D)}\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)&\leqslant\sigma^{2}\|L_{\star}\|_{\rm op}\sqrt{C(\delta)D^{2}\eta^{2}\tau p^{\prime}d^{2}rNT}\,.\\ \end{split}$
@@
Plug in the results from Theorem C.3 and use the definition of $\mathcal{A}(D)\setminus\mathcal{G}(C(\delta),D)$ to lower and upper bound $\|\Delta_{\boldsymbol{A}}\|_{F}$. ∎
Definition C.3.
For applying Theorem B.1 in our setup, consider the following objects:
@@
$1-\exp\left(-C_{HW}C\nu^{2}\eta^{2}\tau p^{\prime}dr\right)\,,$
@@
Applying Theorem B.1 with $P=\tilde{\Delta}_{\boldsymbol{A}}^{\top}\tilde{\Delta}_{\boldsymbol{A}}$ and $r=c^{2}\sigma^{2}\nu\|\tilde{\Delta}_{\boldsymbol{A}}\|_{F}^{2}$,
$\begin{split}&\mathbb{P}\left(\tilde{E}^{\top}\tilde{\Delta}_{\boldsymbol{A}}^{\top}\tilde{\Delta}_{\boldsymbol{A}}\tilde{E}-\mathbb{E}[\tilde{E}^{\top}\tilde{\Delta}_{\boldsymbol{A}}^{\top}\tilde{\Delta}_{\boldsymbol{A}}\tilde{E}]\geqslant c^{2}\sigma^{2}\nu\|\tilde{\Delta}_{\boldsymbol{A}}\|_{F}^{2}\right)\\ &\quad\quad\leqslant\exp\left(-C_{HW}\min{{\left\{\nu^{2}\frac{\|\tilde{\Delta}_{\boldsymbol{A}}\|_{F}^{4}}{\|\tilde{\Delta}_{\boldsymbol{A}}^{\top}\tilde{\Delta}_{\boldsymbol{A}}\|_{F}^{2}},\nu\frac{\|\tilde{\Delta}_{\boldsymbol{A}}\|_{F}^{2}}{\|\tilde{\Delta}_{\boldsymbol{A}}^{\top}\tilde{\Delta}_{\boldsymbol{A}}\|_{\rm op}}\right\}}}\right).\end{split}$
@@
$\begin{split}&\mathbb{P}\left(\|\Delta_{\boldsymbol{A}}E\|_{F}^{2}-\sigma^{2}N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}\geqslant c^{2}\sigma^{2}\nu N\|\Delta_{\boldsymbol{A}}\|_{F}^{2}\right)\leqslant\\ &\quad\quad\exp\left(-C_{HW}\min{{\left\{\nu^{2}N\frac{\|\Delta_{\boldsymbol{A}}\|_{F}^{4}}{\|\Delta_{\boldsymbol{A}}^{\top}\Delta_{\boldsymbol{A}}\|_{F}^{2}},\nu N\frac{\|\Delta_{\boldsymbol{A}}\|_{F}^{2}}{\|\Delta_{\boldsymbol{A}}\|_{\rm op}^{2}}\right\}}}\right).\end{split}$
@@
$\begin{split}1-\exp\left(-C_{HW}C\nu^{2}\eta^{2}\tau p^{\prime}dr\right)\,,\end{split}$
@@
$\operatorname{Tr}(E^{\top}\Delta_{\boldsymbol{A}}L_{\star}E)\leqslant c^{2}\sigma^{2}\nu\|L_{\star}\|_{\rm op}\sqrt{C\eta^{2}\tau p^{\prime}drN}\|\Delta_{\boldsymbol{A}}\|_{F}\,.$
@@
Applying Theorem B.1 with $P=\tilde{\Delta}_{\boldsymbol{A}}$ and $r=c^{2}\sigma^{2}\nu\sqrt{C\eta^{2}\tau p^{\prime}dr}\|\tilde{\Delta}_{\boldsymbol{A}}\|_{F}$,
$\begin{split}&\mathbb{P}\left(\tilde{E}^{\top}\tilde{\Delta}_{\boldsymbol{A}}\tilde{E}-\mathbb{E}[\tilde{E}^{\top}\tilde{\Delta}_{\boldsymbol{A}}\tilde{E}]\geqslant c^{2}\sigma^{2}\nu\sqrt{C\eta^{2}\tau p^{\prime}dr}\|\tilde{\Delta}_{\boldsymbol{A}}\|_{F}\right)\\ &\quad\quad\leqslant\exp\left(-C_{HW}\min{{\left\{\nu^{2}C\eta^{2}\tau p^{\prime}dr,\nu\sqrt{C\eta^{2}\tau p^{\prime}dr}\frac{\|\tilde{\Delta}_{\boldsymbol{A}}\|_{F}}{\|\tilde{\Delta}_{\boldsymbol{A}}\|_{\rm op}}\right\}}}\right).\end{split}$
@@
$\begin{split}&\mathbb{P}\left(\operatorname{Tr}\left(E^{\top}\Delta_{\boldsymbol{A}}E\right)\leqslant c^{2}\sigma^{2}\nu\sqrt{C\eta^{2}\tau p^{\prime}drN}\|\Delta_{\boldsymbol{A}}\|_{F}\right)\geqslant\\ &\qquad\qquad\qquad 1-\exp\left(-C_{HW}\min{{\left\{\nu^{2}C\eta^{2}\tau p^{\prime}dr,\nu\sqrt{C\eta^{2}\tau p^{\prime}drN}\frac{\|\Delta_{\boldsymbol{A}}\|_{F}}{\|\Delta_{\boldsymbol{A}}\|_{\rm op}}\right\}}}\right).\end{split}$
@@
Lemma C.7.
For any $\nu>0$, with probability at least
$1-\exp\left(-C_{HW}\min{{\left\{\nu,\nu^{2}\right\}}}dNT\right),$
@@
Applying Theorem B.1 with $P=I_{dTN},r=c^{2}\sigma^{2}\nu dNT$,
@@
The result follows by the fact $\mathbb{E}[\tilde{E}^{\top}\tilde{E}]=\sigma^{2}dNT$. ∎
Lemma C.8.
@@
$\begin{split}\|\Delta_{{\boldsymbol{A}}_{2}}L_{\star}E\|_{F}^{2}&\geqslant\frac{1}{2}\|\Delta_{{\boldsymbol{A}}_{1}}L_{\star}E\|_{F}^{2}-p^{\prime}\|{\boldsymbol{A}}_{1}-{\boldsymbol{A}}_{2}\|_{\rm op}^{2}\|L_{\star}\|_{\rm op}^{2}\|E\|_{F}^{2}\,.\end{split}$